### v1

# 11email: {giovanni.catalani}[at]airbus.com

Abstract
Recent advances in Neural Fields have enabled powerful, discretization-invariant methods for learning neural operators that approximate solutions of Partial Differential Equations (PDEs) on general geometries. Building on these developments, we introduce enf2enf, an encoder–decoder methodology for predicting steady-state Partial Differential Equations with non-parameterized geometric variability, based on recently proposed Equivariant Neural Field architectures. In enf2enf, input geometries are encoded into latent point cloud embeddings that inherently preserve geometric grounding and capture local phenomena. The resulting representations are then combined with global parameters and directly decoded into continuous output fields, thus efficiently modeling the coupling between geometry and physics. By leveraging the inductive biases of locality and translation invariance, our approach is able to capture fine-scale physical features as well as complex shape variations, thereby enhancing generalization and physical compliance. Extensive experiments on a high-fidelity aerodynamic dataset, a hyper-elastic material benchmark, and multi-element airfoil geometries, demonstrate that the proposed model achieves superior or competitive performance compared to state-of-the-art graph based, operator learning, and neural field methods. Notably, our method supports real time inference and zero-shot super-resolution, enabling efficient training on low-resolution meshes while maintaining high accuracy on full-scale discretizations.
Keywords:
Partial Differential Equations, Neural Operators, Neural Fields
1 Introduction
Partial Differential Equations (PDEs) are fundamental to modeling physical phenomena across numerous scientific and engineering domains [11, 37]. They are essential in fluid dynamics, heat transfer, structural mechanics, electromagnetics, and quantum mechanics mathematical modeling. In aerospace engineering particularly, Computational Fluid Dynamics (CFD) has become an indispensable tool for aircraft design, enabling high-fidelity simulations of complex aerodynamic phenomena. These simulations, primarily based on numerical approximations of the Reynolds Averaged Navier Stokes (RANS) equations [29], provide detailed predictions of flow fields, aerodynamic performance, or structural loads. However, the computational cost of traditional numerical methods remains a significant bottleneck [13]. The complexity arises from the large number of degrees of freedom required to capture relevant physics, mesh refinement requirements for accurate boundary layer resolution, and iterative solution procedures. Additionally, in typical engineering applications, multiple solutions of the same governing equations are required to extensively explore the design space due to variations in domain geometry (boundary conditions) or input conditions, such as fluid density or inflow velocity in the case of aerodynamics. The need for less computationally demanding simulations and the abundance of available data motivate the development of data-driven PDE surrogate models, that can offer remarkable speedups compared to traditional numerical methods, while reducing the level of fidelity [7].
Traditional surrogate models based on modal decomposition such as Proper Orthogonal Decomposition [4] (POD or Principal Component Analysis) coupled with interpolation techniques [31] (Gaussian Process Regression, Radial Basis Functions) excel on fixed geometries and in modeling linear phenomena, but it is not trivial to extend them to handle changing geometries [8] and levels of discretization. Deep neural networks, grounded in universal approximation theory, have demonstrated the ability to represent highly complex functions in high-dimensional spaces, by extracting meaningful representations from complex data [3, 6].
This representational power justifies successes in areas such as image processing [22] and generative modeling, and can extend to the area of PDE modeling, which concerns learning high-dimensional input–output mappings and capturing multiscale phenomena across complex solution manifolds.
Crucially, Deep Learning architectures leverage certain assumptions about the solution space, referred to as inductive biases [16], and restrict the function class to those that align with the underlying dataset or physical constraints, thereby promoting better generalization beyond the training set. Convolutional Neural Networks (CNNs) [21], employ convolution operations enforcing spatial locality and translation invariance. This inductive bias has been shown to be particularly effective for PDE-related tasks [1], where local interactions (e.g., localized regions of fluid flow or stress concentration) are detected in different absolute locations in the spatial domain. The reliance of CNN on uniform grids (or coupling with an interpolation routine) renders them unsuitable for physical applications that employ unstructured meshes [10]. Graph Neural Networks [6] (GNNs) address this issue by extending local message passing to irregular graph topologies, thus accommodating arbitrary mesh structures [28]. However, GNN based models typically encode the mesh topology and resolution directly into their representation, leading to biases when generalizing to unseen discretizations [14] or mesh topologies. Moreover, the requirement to store node neighborhoods and use an increasing number of message passing layers can result in substantial memory and computational overhead, especially for large scale industrial meshes with millions of nodes [18]. These factors hinder generalization across varying mesh resolutions and geometries, emphasizing the need for discretization invariant approaches that can efficiently handle large scale PDEs [25, 9]. Operator Learning aims to directly approximate the PDE solution operator, mapping the input (e.g., boundary conditions, source terms, or geometrical parameters) and output function spaces [2]. By treating both inputs and outputs as functions, it is possible to obtain continuous representations of physical fields that can be queried at arbitrary points in the spatial domain. DeepONet [26] and the Fourier Neural Operator (FNO) [24], are two prominent approaches that have demonstrated compelling results in various PDE modeling tasks. It is not trivial to extend these architectures to handle non-uniformly sampled grids [23] (FFT can only be computed efficiently on structured grids) and unparameterized geometric variations, such as in the case of geometry aware inference, for instance for the predictions of the pressure fields over airfoils.
Conditional Neural Fields have emerged as a powerful framework to construct Neural Operator architectures [38] offering several attractive properties including discretization independence and good spectral convergence thanks to Fourier Feature Networks [36] or sinusoidal [34] layer activations. CORAL [32] encodes latent representations of physical fields (referred to as functa) using a meta-learning approach [12]. For geometry aware inference, two separate encoders are used to encode functa for input functions (geometry) and output physical fields, which are then separately mapped in the latent space: this approach, referred to as encode-process-decode, has also been extended for 3D applications concerning aircraft aerodynamic simulations [9]. Neural Field approaches that rely on a global shared network and latent vector modulation often lack a strong inductive bias in both the neural field architecture and the latent space structure. Because each physical field is encoded as a single global vector, the representation lacks grounding in geometry [39]. This results in latent representations with limited downstream utility, making it challenging to model the effect of local geometric features and large scale physical phenomena[27]. Moreover, existing methods (based on encode-process-decode) typically minimize an error metric on the learned latent space (when mapping input and output latent spaces) rather than directly penalizing discrepancies in the physical output fields themselves, leading to potential mismatches between the latent space objective and the final PDE solution accuracy. Such limitations often necessitate strong regularization on the latent representations to mitigate overfitting [12, 27].
Recent work on Equivariant Neural Fields (ENFs) [39] has shown how incorporating geometric structure in the latent space can significantly improve representation learning and downstream tasks. By encoding fields into latent point clouds that are grounded in the spatial domain, local conditioning and relevant symmetries can be enforced, leading to improved sample efficiency. Locality priors, in particular, are important for PDE modeling (i.e. the state at a given point is predominantly influenced by nearby geometric and flow conditions) and several works have exploited attention mechanisms [33, 40] to implement local conditioning, demonstrating enhanced generalization and reduced computational overhead by focusing the network’s capacity on the most relevant regions of interest.
Accurately modeling geometric variability is critical for predicting physical phenomena: variations of a wing shape or of a structural component directly influence the resulting flow-field or stress distribution, with important consequences for industrial design optimization. Equivariant Neural Fields have been demonstrated for time-dependent PDE modeling on fixed geometries where each time step is encoded into a geometrically grounded latent representation and evolved by a separate latent time-processing module [20]. In contrast, our work addresses a different challenge: modeling variable geometries (and operating conditions) to directly infer the resulting steady-state physical fields. To achieve this, we propose enf2enf, a novel geometry aware operator learning framework that leverages ENFs representations to tackle steady state PDEs on complex geometries. In our encoder-decoder framework, the input geometric fields (such as Signed Distance Functions) are encoded into latent representations that are grounded in the physical domain and are directly mapped to the output fields by a downstream decoder, which handles the processing of both local (geometric) and global (inflow) conditioning. Crucially, compared to previous approaches based on encode-process-decode pipeline [32, 9], we bypass the separate latent space processor which typically introduces an auxiliary latent objective decoupled from the final PDE reconstruction error. This leads to a simpler architecture, reducing computational overhead in training and inference, and yields a more direct interaction between geometry and physics, potentially enhancing generalization performance. Furthermore, we demonstrate competitive performance on a material modeling benchmark and a large scale aerodynamic dataset, characterized by high-resolution CFD simulations over complex airfoil geometries. Our extensive baseline comparisons include state-of-the-art operator learning frameworks, GNN-based methods, and conventional neural field approaches. Detailed analyses reveal that the added inductive biases such as localized geometric encoding and translation equivariant attributes enhance the model’s ability to generalize across varying mesh resolutions and to capture both global inflow conditions and localized phenomena with improved accuracy and physical consistency.
2 Methodology
Encoder Training
Decoder Training
Inference
Figure 1: Architecture overview of enf2enf showing different computational paths during training and inference. The input field is processed by encoder $\mathcal{E}_{\text{a}}$ to produce latent variables $z_{a}$ defined as point clouds (green) with node features $c_{i}^{\text{in}}$, which are nodewise concatenated ($\oplus$) with parameters $\mu$ (leading to output latents $z_{u}$ in red) before being decoded by $\mathcal{D}_{\text{u}}$ to generate the output field.
2.1 Problem Formulation
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded Lipschitz domain, where $d$ is the spatial dimension. We define two separable Banach spaces $\mathcal{A}=\mathcal{A}(\Omega;\mathbb{R}^{d_{a}})$ and $\mathcal{U}=\mathcal{U}(\Omega;\mathbb{R}^{d_{u}})$, consisting of functions mapping from $\Omega$ to $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$ respectively. The input function space $\mathcal{A}(\Omega)$ is defined as:
$\mathcal{A}(\Omega)=\mathcal{A}_{\text{geom}}(\Omega;\mathbb{R})\times\mathcal{A}_{\mu}(\Omega;\mathbb{R}^{d_{\mu}}),$ (1)
where $a_{\text{geom}}\in\mathcal{A}_{\text{geom}}$ encodes the geometry (e.g., Signed Distance Functions continuously describing the shape boundaries), and $\mu\in\mathcal{A}_{\mu}$ represents global parameters (e.g., inflow conditions, material properties). The space $\mathcal{U}$ contains functions mapping to physical field solutions, such as velocity, pressure or stress distribution, that satisfy the governing partial differential equations under the conditions specified by elements of $\mathcal{A}$. The governing partial differential equations define a nonlinear operator mapping between these spaces:
$\mathcal{G}^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}.$ (2)
We seek to approximate $\mathcal{G}^{\dagger}$ with a parameterized operator $\mathcal{G}_{\theta}:\mathcal{A}\rightarrow\mathcal{U}$, where $\theta\in\Theta$ represents the learnable parameterization. Given a probability measure $\nu$ on $\mathcal{A}$, we aim to minimize the approximation error in the continuous setting:
[…]
We seek to learn this operator from a collection of input-output pairs $(a_{i},u_{i})\in\mathcal{A}\times\mathcal{U}$ (where again $a_{i}=(a_{geom,i},\mu_{i})$) such that $u_{i}=\mathcal{G}^{\dagger}(a_{i})$. Due to the numerical nature of the solutions, at training time we have access to $n_{tr}$ pairs of input and output functions $(a_{i},u_{i})_{i=1}^{n_{tr}}$ evaluated on unstructured grids. Specifically, each sample is observed on its own computational mesh $\mathcal{X}_{i}$. We denote by $a|_{\mathcal{X}_{i}}=(a(x))_{x\in\mathcal{X}_{i}}$ and $u|_{\mathcal{X}_{i}}=(u(x))_{x\in\mathcal{X}_{i}}$ the vectors of function values evaluated over the sample mesh. This leads to the empirical minimization problem:
$\theta^{\dagger}=\arg\min_{\theta\in\Theta}\frac{1}{n_{tr}}\sum_{i=1}^{n_{tr}}\sum_{x\in\mathcal{X}_{i}}\|u_{i}(x)-\mathcal{G}_{\theta}(a_{i})(x)\|^{2}.$ (4)
2.2 Model Architecture and Training
Our architecture uses two main components: $f_{\theta_{a},z_{a}}$, $f_{\theta_{u},z_{u}}$ for input and output functions respectively, both implemented as Equivariant Neural Fields (ENFs) [39], where $\theta_{a},\theta_{u}$ are global shared parameters and $z_{a},z_{u}$ are sample-specific latent point cloud representations of input and output functions. Input functions $a_{i}$ are encoded into latent representations using the input encoder $z_{a_{i}}=\mathcal{E}_{\text{a}}({a_{i}})$ by solving an optimization problem $f_{\theta_{a},z_{a_{i}}}(x)|_{\mathcal{X}_{i}}=a_{i}$. This procedure, known as auto-decoding, is better detailed in the following dedicated paragraph. Obtained input representations are then (nodewise) concatenated with global parameters $\mu$ (when present), and decoded to output physical fields through the output spatial decoder $\mathcal{D}_{u}$:
$\hat{u}_{i}(x)=D_{u}(z_{u_{i}})(x)=f_{\theta_{u},z_{u_{i}}}(x),\quad z_{u_{i}}=\text{cat}\big{(}z_{a_{i}},\,\mu\big{)}.$ (5)
The resulting operator $\mathcal{G}_{\theta}$ can then be expressed as the composition:
$\mathcal{G}_{\theta}=\mathcal{D}_{\text{u}}\circ\mathcal{E}_{\text{a}}.$ (6)
Training proceeds in two stages, sketched in Figure 1, by minimizing the reconstruction losses for both input and output fields:
$\mathcal{L}^{a}=\sum_{i=1}^{n_{tr}}\|a_{i}-\mathcal{D}_{\text{a}}(\mathcal{E}_{\text{a}}(a_{i}))\|^{2}_{\mathcal{A}},\quad\mathcal{L}^{u}=\sum_{i=1}^{n_{tr}}\|u_{i}-\mathcal{D}_{\text{u}}(\mathcal{E}_{\text{a}}(a_{i}))\|^{2}_{\mathcal{U}}$ (7)
where the input decoder component $D_{a}$ is used for learning to reconstruct the input functions from input latents, and it is not involved in the inference process. In the following we describe the specific architecture of input and output ENFs, and the encoding process.
2.2.1 Translation Equivariant Neural Fields.
Equivariant Neural Fields (ENFs) are coordinate-based neural networks that can be equipped with different symmetries through the choice of bi-invariant attributes [39]. These attributes define the relationship between spatial coordinates and latent representations in a way that preserves the desired symmetries and invariances of the problem. Compared to standard conditional Neural Fields autodecoder architectures that extract sample-wise representations in the form of 1D arrays, ENFs encode functions on latent point cloud representations embedded in the physical domain and as such grounded in geometry [20]:
$z:=\{(p_{i},c_{i})\}_{i=1,..,N_{lat}}$ (8)
where the latent features $c_{i}\in\mathbb{R}^{l_{d}}$ are localized in the latent positions $p_{i}\in\mathbb{R}^{d}$ of the physical domain. For fluid flow problems and general physical PDEs without evident rotational or scaling symmetries, translation equivariance represents a strong inductive bias to characterize the physical system. Many fundamental physical phenomena, such as vortices or shocks, can occur at different absolute locations while maintaining identical local behavior and properties. Therefore, we employ relative positions as our bi-invariant attribute:
$b_{i,m}=\gamma(x_{m}-p_{i})$ (9)
where $p_{i}$ represents the position of the $i$-th latent point and $x_{m}$ is the query position. Fourier Feature Encoding $\gamma$ is used on the inputs, as in standard Neural Fields to overcome the spectral bias and capture the high frequency components of the target signals. The output of the neural field architecture at a specific location $x_{m}$ is obtained as the weighted sum of the different contributions from the latent points:
$f_{\theta_{a}}(x_{m};z)=W_{o}\sum_{i=1}^{N_{lat}}\text{att}_{m,i}v(b_{i,m},c_{i})$ (10)
where the attention weights are given by:
$\text{att}_{m,:}=\text{softmax}_{i=1,...,N_{lat}}\left(\frac{q(b_{i,m})^{T}k(c_{i})}{\sqrt{d_{k}}}-\sigma||x_{m}-p_{i}||^{2}\right)$ (11)
and the second term (whose magnitude is controlled by the gaussian window size scalar $\sigma$) penalizes the distance of the query point from the latent position, therefore enhancing local conditioning. The query and key functions are parameterized as:
$q(b_{i,m})$ $=W_{q}b_{i,m}$ (12)
$k(c_{i})$ $=W_{k}c_{i}$ (13)
while the value is obtained by linear transformation of the bi-invariants through scale and shift modulation:
$v(b_{i,m},c_{i})=(W_{v}c_{i})\odot(W_{a}^{\gamma}b_{i,m})+W_{a}^{\beta}b_{i,m}.$ (14)
2.2.2 Encoding
For each input function $a$ (geometric field), the encoding process follows a meta-learning optimization procedure based on CAVIA [41]. Specifically, given the shared network parameters $\theta_{a}$, the latent point-cloud representations $z_{a}=\{(p_{i},c_{i})\}_{i=1,...,N_{lat}}$ are obtained by solving an optimization problem (omitting the index $i$ for brevity):
$c^{(0)}$ $=0\quad p^{(0)}=p^{(K)}$ (15)
$c^{(k+1)}$ $=c^{(k)}-\alpha\nabla_{c^{(k)}}\mathcal{L}^{a}_{\mathcal{X}_{i}}(f_{\theta_{a},c^{(k)}},a)\quad\text{for }0\leq k\leq K-1.$ (16)
where compared to [39, 20], we initialize the latent positions $p$ for all samples to cover uniformly the input domain and decide not to optimize them, as the updating their value results in less stable training results. Here, $\alpha$ is the inner loop learning rate, $K$ is the number of optimization steps (typically set to 3), and $\mathcal{L}^{a}_{\mathcal{X}_{i}}$ is the reconstruction loss evaluated over the training grid defined on the input domain. During training of the input encoder, the optimization problem (inner loop) is solved in parallel with the global parameters $\theta_{a}$ optimization (outer loop). At inference, the global parameters are fixed, as it is sufficient to run the inner loop in few gradient descent steps ($K$) to encode a new geometry.
3 Experiments
We experimentally demonstrate the validity of the proposed approach on two steady state PDE benchmarks: (i) the hyper-elastic material dataset and (ii) the AirFRANS Dataset, consisting of numerical solutions of the incompressible Navier-Stokes equations on 2D airfoils. Additionally, we provide an experiment on learning more complex geometry representations of multi-element airfoils.The implementations and the code to reproduce the experiments can be found in the dedicated repository: https://anonymous.4open.science/r/enf2enf-322F.
3.1 The Hyper-elastic Material Dataset
3.1.1 Description
We consider a hyper-elastic material benchmark problem commonly used for geometry aware inference [23]. The boundary-value problem is defined on a unit cell domain, featuring an arbitrarily shaped void at its center. The underlying finite element solver employs approximately 100 quadrilateral elements. The input configuration is provided as a point cloud of around 1000 points representing the geometry, and the target output is the corresponding stress field. The dataset comprises 1000 training samples and 200 test samples. Although not representing a large scale PDE benchmark, this is a good demonstrative example that the proposed model has competitive performance beyond fluid dynamics applications. More details about the dataset are given in the Supplementary Materials.
3.1.2 Experimental Details
For the hyper-elastic material dataset, our approach employs an encoder-decoder architecture without concatenating any global features. For each input function, we compute a geometry descriptor by encoding the displacement field between the average mesh position and the current sample’s points, as done in previous works [32]. The obtained latent representation is defined on a point cloud of 9 nodes with 8-dimensional features, which is subsequently used to condition the output decoder model for the stress field. Baseline models such as FNO and UNet are limited to uniform grids: these methods rely on interpolating the unstructured input data onto a uniform grid, processing the data, and then re-interpolating the output back onto the original mesh. We also compare against DeepONet [26], functa2functa (CORAL with random Fourier features), and CORAL (which employs SIREN activation functions).
3.1.3 Results
Table 1 reports the L2 mean relative errors on the hyper-elastic material dataset. The results indicate that the performance of FNO and UNet is limited, likely due to the interpolation steps required to map between unstructured and uniform grids. Both CORAL and functa2functa achieve lower errors, with CORAL performing best among the baselines. Although the difference between local (enf2enf) and global latent encoding (functa2functa) is less pronounced for this relatively simpler PDE where multiscale information is not as critical, the proposed enf2enf still yields competitive results.
Table 1: Comparison of Different Models. We show results on the AirFRANS Dataset (left) with Mean Squared Error (MSE) on the normalized predictions of the pressure field on the volume and on the surface, as well as the MSE error and the Spearman Correlation Coefficient $\rho_{L}$ on the lift coefficient predictions; and on the Hyper-elastic Material Dataset (right) with Mean L2 Relative Errors on the test set. Bold indicates best, underlined indicates second best.
Model AirFRANS
Volume $\downarrow$ Surface $\downarrow$ $C_{L}$downarrow $\rho_{L}$uparrow
Simple MLP 0.0081 0.0200 0.2108 0.9932
GraphSAGE 0.0087 0.0184 0.1476 0.9964
MeshGraphNet 0.0214 0.0387 0.2252 0.9945
Graph U-Net 0.0076 0.0144 0.1677 0.9949
GINO 0.0297 0.0482 0.1821 0.9958
Transolver 0.0037 0.0142 0.1030 0.9978
functa2functa* 0.0045 0.0140 0.1591 0.9964
enf2enf (Ours) 0.0026 0.0060 0.0761 0.9983
Model Hyper-elastic
Test Error $\downarrow$
FNO 4.95e-2
UNet 5.34e-2
GraphNO 1.27e-1
DeepOnet 9.65e-2
Geo-FNO 3.41e-2
CORAL 1.67e-2
functa2functa* 2.19e-2
enf2enf (Ours) 1.88e-2
3.2 The AirFRANS dataset
3.2.1 Description
The AirFRANS dataset [5] provides a challenging benchmark for surrogate modeling in computational fluid dynamics, simulating the incompressible Reynolds-averaged Navier-Stokes (RANS) equations over a wide range of airfoil geometries and flow conditions. The geometries are systematically obtained through variations of the NACA 4- and 5-digit series airfoils, resulting in a diverse design space. The implicit distance (or SDF) field is defined over the entire mesh for each simulation as a continuous descriptor of the airfoil shapes. The flow conditions cover incompressible regimes (Mach numbers < 0.3) with angles of attack ranging from -5° to 15°, representing typical subsonic flight conditions. The Reynolds number, ranging from 2 to 6 Millions is set within realistic flight ranges (for relatively small low speed aircrafts). The simulations are performed using OpenFOAM [19]. The dataset features a large scale PDE problem of industrial relevance, discretized on high resolution C-grid meshes with approximately 200,000 points per geometry, refined towards the airfoil surface to capture boundary layer effects. Each simulation takes around 25 minutes to achieve convergence using 16 CPU cores of an AMD Ryzen Processor. We focus on the full task as 800 samples are used for training and 200 test samples for model evaluations. More details about the Airfrans Dataset are given in the Supplementary Materials.
3.2.2 Experimental Setting and Baselines
We mainly evaluate the model performance on the volumetric and surfacic pressure field predictions and the resulting lift coefficients, as these are the primary quantities of interest for aerodynamic design. Our methodology (enf2enf) relies on a two step training strategy: (i) Encoding, the encoder ENF takes as input the coordinates ($x,y$) and fits the implicit distance fields using latent point clouds (9 points with 8-dimensional latent features). (ii) Decoding, the fitted input latent point clouds are concatenated with the freestream conditions ($V_{x},V_{y}$) and decoded to the output fields by a similar ENF architecture which operates in decode-only mode (latent features are not updated during training). More details about implementation and hyperparameters are provided in Supplementary Materials.
In order to show the effect of the added inductive bias on accuracy and generalization, we compare our model to a strong Neural Field baseline based on meta-learned functa [12] (referred to as functa2functa). This implements the CORAL [32], but using Fourier Features rather than SIREN networks used in the original paper (this would otherwise lead to poor convergence on this dataset). Additional comparison are performed with state-of-the-art architectures based on: Graph Neural Networks (including GraphSAGE [17], MeshGraphNet [28], and Graph U-Net [15]), Neural Operators (GINO [25]), and Transformer-based models (Transolver [40]).
3.2.3 Results
In Table 1, it can be observed that enf2enf achieves the lowest errors across volumetric and surface pressure predictions as well as the lift coefficient estimation. The performance gap becomes more evident especially on the surface pressure, as more complex flow features are localized close to the boundary, compared to the rest of the domain where the flow can be considered quasi-potential. Graph Neural Network (GNN) baselines—such as GraphSAGE and Graph U-Net tend to perform better than MeshGraphNet for large scale PDEs thanks to the use of pooling and aggregation operations over extensive neighborhoods. Similarly to Graph U-Net, the proposed method, performs feature encoding on a latent set of nodes. However, Graph U-Net relies on complex pooling operations and aggregation steps to produce its latent representation, which is ultimately determined by the discretization of the input graph. Instead, enf2enf as a neural operator approach, yields representations that are independent of the specific input mesh instance, and can be decoded continuously in the spatial domain, thus being better suited to handle such tasks. To assess the impact of explicit local feature encoding on capturing relevant aerodynamic phenomena, we next compare with global latent encoding methods such as functa2functa in Figure 2. This aspect is especially crucial in fluid dynamics applications, where localized phenomena—such as the leading edge suction peak—play a critical role in determining aerodynamic performance. As depicted in Figure 2, the surface $C_{p}$ plot reveals that the maximum error at the leading edge is mitigated in the enf2enf predictions. Moreover, the isobaric lines outputs match more closely the high fidelity simulation with lower levels of noise (overfitting) and improved physical compliance. While Transolver employs a similar inductive bias to capture local effects, it lacks a more explicit mechanism for encoding geometric information, which may account for its relatively inferior performance. Finally, the t-SNE visualizations in Figure 3 further illustrate that enf2enf effectively learns and disentangles local geometric features, yielding a well structured latent space. The local latent codes computed at farfield and central nodes (left and center panels) exhibit clear clustering that correlates with key geometric parameters, such as airfoil thickness. Notably, the farfield latent feature (left panel) shows little sensitivity to thickness variations because its spatial position is distant from the airfoil, and therefore it does not contribute significantly to reconstructing the SDF field in regions where thickness is a critical factor. In contrast, the 64-dimensional array latent representation from functa2functa (right panel) is less structured, as it aggregates local features into a single global representation. Consequently, the localized encoding in enf2enf enables the decoder to focus on significant aerodynamic features while filtering out less relevant local variations.
Furhtermore, we evaluate discretization convergence on the Airfrans dataset by training each model at lower mesh resolutions (5k, 10k, 32k) and then testing on the full-scale meshes consisting of approximately 200k points. For direct inference, the entire high resolution mesh is processed in a single forward pass. For iterative inference, the mesh is divided into multiple subsamples, each processed in a separate forward pass until all 200k nodes are covered. Our results show that graph based methods (GraphSAGE and GUNet) perform poorly under direct inference when the test resolution greatly exceeds the training resolution, although iterative inference can partially mitigate this at the cost of multiple runs. For instance, GraphSAGE at 32k training resolution requires roughly six passes to cover a 200k test mesh, significantly increasing total inference time. By contrast, Transolver can handle zero-shot super-resolution training at 5k and testing at 200k yet sees degraded accuracy when the resolution gap grows. enf2enf remains discretization invariant and shows no difference between direct and iterative inference; it achieves high accuracy even when trained at much lower resolutions, thus avoiding the cost of large scale training while maintaining excellent performance at full resolution. All models achieve a remarkable 4 to 5 order of magnitude speedup at inference compared to high fidelity CFD. More results and visualization are provided in the Supplementary Material section.
Figure 2: Visualization of pressure field predictions on an airfoil geometry. Top row: predictions from enf2enf (left) and functa2functa (middle) compared to the ground truth (right). Bottom row: corresponding absolute errors for each model and surface pressure distribution.
enf far-field                       enf close                          functa
Figure 3: t-SNE visualizations of latents fitted on the SDFs of AirFrans test split. Left: 8-dimensional latent code computed at a farfield point of the latent point cloud. Center: latent code at a central node located on the airfoil. Right: t-SNE visualization of the 64-dimensional global embedding obtained via functa.
Figure 4: Discretization convergence analysis on the AirFRANS dataset. Left: Direct inference performance vs training resolution Center: Iterative inference; note that ENF is omitted in this panel due to its coordinate-based design. Right: Inference time histogram for models trained at a resolution of 32k points.
3.3 Additional Experiment: Multi-element Airfoil
To further demonstrate the benefits of localized geometric encoding, we conducted additional experiments on shape embedding of multi-element airfoils, by generating a synthetic dataset consisting of 200 geometric variations derived from the NASA GA(W)-I airfoil with a fowler flap. The variations are created through systematic translation and rotation of the flap element while maintaining the main airfoil fixed, representing typical high-lift device operations in aircraft design. For each configuration, we computed the signed distance function (SDF) field and trained both enf and functa encoders for reconstruction accuracy, evaluating the models’ capacity to learn structured representations of multi-component geometries with relative positioning and orientation variations. Dramatic improvement in reconstruction accuracy (4 orders of magnitude) are obtained using the ENF encoder: while functa encoder treats each geometric variation as an entirely new shape to be encoded in its global latent space, ENF translation-equivariant architecture naturally handles the relative positioning of elements through its local feature encoding. This allows the model to effectively decompose the geometry into its constituent elements and capture their spatial relationships. The qualitative results in Figure 5 display how the ENF-based encoder accurately reconstructs both the main airfoil and flap geometries across different configurations, maintaining sharp features and precise relative positioning.
Figure 5: Shape reconstruction results on the multi-element airfoil geometry. For functa latent of shape 32 are used. For enf latents point clouds of 4 nodes and 8-dimensional node features are used. First row: high-lift configuration. Second row: high-lift and high-drag configuration.
4 Discussion and Limitations
Respecting the underlying physics is paramount in PDE modeling. Although approaches that incorporate explicit physics constraints in the loss function, such as Physics Informed Neural Networks (PINNs) [30], have shown promise, they remain limited by challenges in generalizing to complex shape variations, accommodating varying boundary conditions, and scaling beyond simple test cases. By contrast, restricting the search space to functions that inherently respect the underlying symmetries of the PDE operator provides an alternative pathway for incorporating physical constraints. In this work we propose a framework to learn steady state PDE surrogates using the recently introduced Equivariant Neural Fields [39], combining the geometric inductive bias typical of CNNs and GNNs with the discretization invariance of neural fields. These properties are beneficial for large scale PDE modeling, where training on full resolution meshes is computationally prohibitive, and for capturing the effects of geometric variability. Nevertheless, several limitations remain. First, the positions of the latent points are currently fixed a priori. Preliminary experiments aimed at learning these positions resulted in increased computational cost and training instability. Future work may explore adaptive strategies to allow latent points to self organize around more relevant spatial regions, thereby potentially enhancing the model’s capacity to capture critical geometric features. Additionally, while enf2enf has been demonstrated on two dimensional problems, its extension to three-dimensional applications remains to be investigated. In 3D, a larger number of latent points will likely be necessary to adequately cover the domain, which may introduce further challenges in terms of model complexity and computational efficiency.
5 Conclusion
In this work, we introduced a novel geometry aware operator learning framework that leverages Equivariant Neural Fields (ENFs) within an encoder–decoder architecture to address steady state PDEs on complex geometries. The method encodes the input geometry using localized latent embeddings and directly decodes these representations into continuous output fields, thereby achieving discretization invariance. Extensive experiments on a large scale aerodynamic dataset (AirFRANS) and a hyper-elastic material benchmark demonstrate that our approach attains competitive and in many cases superior performance compared to state-of-the-art baselines, including graph based methods, transformer models, and alternative neural field techniques.
{credits}
[…]
The authors have no competing interests.
Appendix 0.A Implementation Details
The implementation of our model architecture, training scripts, and dataset processing pipelines is available at https://anonymous.4open.science/r/enf2enf-322F. This repository contains all the code necessary to reproduce our experiments. Our proposed enf2enf model is implemented in JAX, building upon the original Equivariant Neural Fields codebase. For the baseline, we generally use the existing codebases in Pytorch.
0.A.1 enf2enf
The enf2enf architecture consists of two main components: an encoder network ($f_{\theta_{a}}$) that learns latent representations of input geometries, and a decoder network ($f_{\theta_{u}}$) that maps these representations to physical fields. Both components utilize Equivariant Neural Fields (ENFs) with translation equivariance as the core building block. The encoder follows a meta-learning optimization procedure based on CAVIA [41], where for each input function, the latent features $c_{i}$ are optimized through $K$ inner steps with learning rate $\lambda_{c}$, while the network parameters $\theta_{a}$ are updated through standard gradient descent. Each ENF employs a multi-head attention mechanism where query points attend to a set of latent points $z:={(p_{i},c_{i})}_{i=1,..,N_{lat}}$. The attention scores are computed using both positional information and feature embeddings, with a Gaussian window (of size $\sigma$) function promoting locality. Random Fourier Features (RFF) are used to encode coordinate information, where inputs are projected using fourier features following $\gamma(x)=[\cos(Wx),\sin(Wx)]$ where $W\sim\mathcal{N}(0,\sigma_{RFF}^{2})^{d\times d_{in}}$. The latent positions $p_{i}$ are initialized to uniformly cover the input domain and remain fixed during training to ensure stability.
Table 2: Model Architecture and Training Specifications
Hyper-parameter AirFRANS Elasticity
$f_{\theta_{\text{a}}}$ / $f_{\theta_{\text{u}}}$ Learnable Parameters $\theta_{\text{a}},\{c_{i}\}_{i=1}^{n_{\text{lat}}}$ / $\theta_{\text{u}}$ $\theta_{\text{a}},\{c_{i}\}_{i=1}^{n_{\text{lat}}}$ / $\theta_{\text{u}}$
Gaussian Window $\sigma$ 0.1 0.1
Attention Heads 2 2
RFF $(d,\sigma_{RFF})$ $(128,1)$ / $(256,2)$ $(128,2)$ / $(256,10)$
Latent Dims $(n_{\text{lat}},l_{\text{dim}})$ $(9,8)$ / $(9,8+2)$ $(9,8)$ / $(9,8)$
Network Width 128 / 256 128 / 256
N. Attention Blocks 0 / 2 0 / 2
Optimization LR $\lambda_{\theta}$ 1e-4 / 1e-4 1e-4 / 1e-4
Inner Steps $K$ 3 / - 3 / -
Inner LR $\lambda_{c}$ 1.0 / - 0.5 / -
Downsampling 10k / 10k Full mesh
Epochs 800 / 2500 1500 / 2500
0.A.1.1 AirFRANS
For the AirFRANS dataset, we initialize the latent positions in a bounding box $[-0.5,0.25]\times[-0.25,0.25]$ closer to the airfoil surface, as most of the interesting flow phenomena occur in this region. The latent feature dimension in the decoder is increased from 8 to 10 to accommodate the concatenation of inflow conditions (velocity components) that globally condition the output field generation. The dataset is normalized by standardizing the output pressure fields and scaling the input coordinates to $[-1,1]$ through min-max normalization. Following the ’full’ task setup, we use 800 samples for training and 200 for testing. Due to the large mesh sizes ($\sim 200,000$ points), we dynamically downsample to 10,000 points during both encoder and decoder training. At inference time, the continuous nature of Neural Fields allows us to evaluate on the full resolution meshes.
0.A.1.2 Elasticity
For the elasticity problem, latent positions are initialized in a larger bounding box $[-0.75,0.75]\times[-0.75,0.75]$ to better cover the unit cell domain. The inputs and outputs are normalized following the same strategy as AirFRANS. Given the relatively small mesh sizes ($\sim 1000$ points), we can process the entire point cloud during training without downsampling, leading to better reconstruction of sharp features near the void boundaries.
0.A.2 Baseline Models
We compare our approach against several state-of-the-art methodologies across different architectural paradigms:
0.A.2.1 Neural Field Baselines
We implement functa2functa, based on the CORAL architecture but modified to use Random Fourier Feature encoding instead of the original SIREN networks. To ensure a fair comparison, we employ a 64-dimensional latent space, which approximately matches our model’s representational capacity (9 latent points × 8 features). The Fourier Feature encoding parameters are kept consistent with our approach to isolate the impact of the architectural differences.
0.A.2.2 Graph-Based Approaches
We evaluate against the graph neural network baselines presented in the AirFRANS benchmark, including GraphSAGE, MeshGraphNet, and Graph U-Net. For computational efficiency on the AirFRANS dataset, these models operate on downsampled meshes of 32,000 points while maintaining the original mesh connectivity structure. This sampling strategy provides a balance between computational feasibility and preservation of important flow features.
0.A.2.3 Neural Operator Methods
For the AirFRANS dataset, we compare against Transolver, maintaining their original implementation settings with input meshes downsampled to 32,000 points. On the elasticity benchmark, we include comparisons with Geo-FNO and traditional convolutional architectures (U-Net, FNO) as reported in previous studies. It’s worth noting that the grid-based methods (U-Net, FNO) require an additional interpolation step to map between the unstructured mesh and regular grid representations, which can impact their performance on complex geometries.
0.A.3 Hardware and Runtime
All experiments were performed on NVIDIA A100 GPUs with 40GB of memory. Training times and epochs for different models (only for the models retrained on our hardware) are reported in Table 3.
Table 3: Training Time and Epochs Across Different Models and Datasets
AirFRANS Elasticity
Model Time (hrs) Epochs Time (hrs) Epochs
enf2enf (encoder/decoder) 5.7/12.9 800/2500 4 1500
functa2functa (encoder/process/decoder) 15 2000 10 2000
GraphSAGE 5.8 400 3 1000
Graph U-Net 8.5 400 5 1000
Transolver 20.7 400 - -
Training times include all necessary pre-processing steps, such as mesh downsampling for graph-based methods and interpolation for grid-based approaches. For enf2enf and functa2functa, we report encoder and decoder training times separately as they are trained sequentially.
Appendix 0.B Datasets Details
0.B.1 AirFRANS Dataset
The AirFRANS dataset consists of numerical solutions of the Reynolds-Averaged Navier-Stokes (RANS) [29] equations for incompressible flow around airfoil geometries [5]. The RANS equations are derived by applying Reynolds decomposition to the velocity and pressure fields:
$u_{i}=\bar{u}_{i}+u_{i}^{\prime},\quad p=\bar{p}+p^{\prime}$ (17)
where $\bar{\cdot}$ denotes the ensemble-averaged quantity and $\cdot^{\prime}$ represents the fluctuating component. Substituting these decompositions into the incompressible Navier-Stokes equations and taking the ensemble average leads to:
$\partial_{i}\bar{u}_{i}=0$ (18)
$\partial_{t}\bar{u}_{i}+\partial_{j}(\bar{u}_{i}\bar{u}_{j})=-\frac{1}{\rho}\partial_{i}\bar{p}+\nu\partial^{2}_{jj}\bar{u}_{i}-\partial_{j}(\overline{u^{\prime}_{i}u^{\prime}_{j}})$ (19)
The Reynolds stress tensor $\tau_{ij}=-\overline{u^{\prime}_{i}u^{\prime}_{j}}$ is modeled using the Boussinesq hypothesis, which introduces the concept of turbulent viscosity $\nu_{t}$:
$\tau_{ij}=\nu_{t}\left(\partial_{i}\bar{u}_{j}+\partial_{j}\bar{u}_{i}\right)-\frac{2}{3}k\delta_{ij}$ (20)
where $k=\frac{1}{2}\overline{u^{\prime}_{i}u^{\prime}_{i}}$ is the turbulent kinetic energy. Incorporating this model and defining an effective pressure $\bar{p}:=p+\frac{2}{3}\rho k$, the final form of the incompressible RANS equations becomes:
$\partial_{i}\bar{u}_{i}=0$ (21)
$\partial_{t}\bar{u}_{i}+\partial_{j}(\bar{u}_{i}\bar{u}_{j})=-\frac{1}{\rho}\partial_{i}\bar{p}+\partial_{j}[(\nu+\nu_{t})\partial_{j}\bar{u}_{i}]$ (22)
In the AirFRANS dataset, these equations are solved using the Spalart-Allmaras (SA) turbulence model [35], which introduces a single transport equation for a modified eddy viscosity $\tilde{\nu}$. The turbulent viscosity is then computed as $\nu_{t}=f_{v1}\tilde{\nu}$, where $f_{v1}$ is a damping function that ensures proper near-wall behavior. The SA model is particularly well-suited for aerodynamic applications as it was specifically calibrated for aerodynamic flows with adverse pressure gradients and separation. The dataset provides numerical solutions obtained using OpenFOAM’s [19] incompressible solver with the SA turbulence model. The simulations are performed on high-resolution C-grid meshes ($\sim$200,000 cells) with appropriate refinement near the airfoil surface to resolve the boundary layer ($y^{+}<1$). The flow conditions span Reynolds numbers from 2 to 6 million, with angles of attack ranging from -5° to 15°. The entire dataset generation process, validation and benchmarking is described in the original paper [5].
0.B.2 Hyper-elastic Material Dataset
The structural mechanics problem considered in this work involves the analysis of a hyperelastic material described by the momentum conservation equation:
$\rho_{s}\frac{\partial^{2}u}{\partial t^{2}}+\nabla\cdot\sigma=0$ (23)
where $\rho_{s}$ represents the material density, $u$ denotes the displacement field, and $\sigma$ is the Cauchy stress tensor. The problem domain consists of a unit square $\Omega=[0,1]\times[0,1]$ containing a centrally located void with variable geometry. To generate physically meaningful void configurations, a stochastic process for the void radius is defined as:
$r=0.2+\frac{0.2}{1+\exp(\tilde{r})},\quad\tilde{r}\sim\mathcal{N}(0,4^{2}(-\nabla+32)^{-1})$ (24)
This formulation naturally enforces the physical constraint $0.2\leq r\leq 0.4$ on the void radius. The material behavior is characterized by an incompressible Rivlin-Saunders constitutive model, setting the material parameters to $C_{1}=1.863\times 10^{5}$ and $C_{2}=9.79\times 10^{3}$, representing a soft rubber-like material. The boundary conditions consist of: fixed displacement (clamped) condition on the bottom edge, traction force $t=[0,100]$ and traction-free conditions on all other boundaries. The dataset was generated using a finite element solver employing approximately 100 quadratic quadrilateral elements, with each simulation requiring around 5 CPU seconds to complete. The input geometry is represented by point clouds containing roughly 1000 points that describe the void shape and domain boundaries. The target output fields consist of the stress components computed at these same locations. The dataset split used in this paper experiments comprises 1000 training samples and 200 test cases, following the settings of the orginal work introducing the dataset [23].
Appendix 0.C Additional Results
Figure 6: Contour plots from the elasticity dataset. Predictions for the stress field and target values with the corresponding absolute errors.
Figure 7: Violin plot of the distribution of relative errors on $C_{L}$ for ENF and functa on the AirFRANS test dataset. The plot includes markers indicating the mean and median of the error distributions. Individual data points are randomly jittered along the x-axis to aid visualization.